Given functions f and g , perform the indicated property.
f(x)=9x-2, g(x)=2x-3
Find fg.

Sorry didnt know where to post it again.

Answer by RedemptiveMath(80)

This problem expresses two functions, f(x) and g(x). It wants the product of these two, or fg, (f*g)(x) or f(x)*g(x). All you need to do is multiply what f(x) equals and what g(x) equals. So,

f(x)*g(x) =
(9x-2)(2x-3) =
18x^2 - 27x - 4x + 6 =
18x^2 - 31x + 6.